# The Core of Approval-Based Committee Elections with Few Candidates

### v1

@@ 4.3 Committee size $k\geqslant 9$
The PAV-based technique that worked for up to $k=8$ does not continue to work for $k=9$, since there are examples where PAV selects a unique committee which fails to be in the core. The following example has this property, and it is the smallest such example with respect to the number of voters ($n=27$).
@@ 5 Few Candidates
Inspecting the examples in Section 4 where PAV fails the core, we see that they are well-structured. Indeed, they are even laminar instances in the sense of (Peters and Skowron, 2020, Definition 2), and it is easy to see that on these profiles, a core-stable committee does exist. Thus, there is some hope to prove existence of core-stable committees by “patching” the PAV committee when it fails to be in the core.
We will define an artificial rule, based on PAV, that we will show satisfies core stability for up to $m=15$ candidates. On a high level, the rule first computes a PAV committee, and checks if it satisfies the core. If so, it returns it. If not, and $T$ is a deviation from $W$, it then deletes all voters who prefer $T$ to $W$, and computes a PAV committee with respect to the remaining voters, but subject to the constraint that $T\subseteq W$. It then checks if the result is in the core; if not, it adds additional constraints until it reaches a core-stable committee. This rule is formally described using pseudocode in Algorithm 1.
@@ 5.1 Analysis of the Method
where $\mathcal{A}_{t}=\{A\in\mathcal{A}:u_{A}(T_{s})\leqslant u_{A}(W_{s})\text{ for}s=1,\dots,t-1\}$ is the set of “active” ballots. That is, $W_{t}$ locally maximizes the PAV score among all committees that include all prior deviations, taking only those voters into account that did not participate in prior deviations.
@@
              if $(W_{t},T_{t})$ is not canonical then
@@
Thus, to prove the existence of core-stable committees, it suffices to enumerate all histories and check that they fulfil the condition of Proposition 5.2. Given a potential history, one can check using an LP solver whether it is a history by checking whether the system of linear inequalities in Definition 5.1 has a solution. This way, we can compute the set of histories using a standard breadth-first search, as shown in Algorithm 2. A key insight is that we only need to consider “canonical” histories in our enumeration. For example, we may assume without loss of generality that $W_{1}$, the first committee of the history, is $\{c_{1},\dots,c_{k}\}$. Similarly, we do not need to consider all potential deviations $T_{1}$, as it suffices to take one for each possible combination of the size $|T_{1}\cap W_{1}|$ and of $|T_{1}\setminus W_{1}|$. Similar symmetry-breaking conditions apply for later steps.
@@
The recursive PAV rule fails for $m=16$, $k\in\{10,11\}$. For $m=16$, $k=10$, the smallest failure example I have found has 40 448 550 voters (though smaller ones surely exist). The example is available online. [^5] This example implements the history ($(\{c_{0},c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{9}\},\{c_{0},c_{10},c_{11}\}),$ ($(\{c_{0},c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{10},c_{11},c_{12}\},\{c_{13},c_{14},c_{15}\}),$ ($(\{c_{0},c_{1},c_{2},c_{3},c_{10},c_{11},c_{12},c_{13},c_{14},c_{15}\},\{c_{4},c_{5},c_{6},c_{7},c_{8}\})$. The recursive PAV rule does work for $m=16$, $k\in\{9,12,13,14\}$, and it is plausible that it can be fixed ad hoc for $k\in\{10,11\}$, so it is likely that the core continues to exist for $m=16$.
@@ 6 Droop Quota
On this profile, $W=\{c_{1},c_{2},c_{5},c_{6},c_{7},c_{8}\}$ is the unique PAV committee. However $W$ is not in the Droop core: consider $T=\{c_{1},c_{2},c_{3},c_{4}\}$, which has support from $\frac{14}{24}\approx 0.583$ of the voters, while $|T|/(k+1)=\frac{4}{7}\approx 0.571$ is strictly smaller. ∎
@@ 7 Conclusions
Our approach was based on linear programming, and in particular this approach allowed us to reason independently of the number of voters. The PAV rule and its variants are particularly well-suited for these LP formulations. However, finding core counterexamples for many other rules is not possible using similar linear programs. For example, the Method of Equal Shares (MES) (Peters and Skowron, 2020) or the sequential Phragmén method (Phragmén, 1894; Janson, 2016) seem not to admit linear formulations (because they would require multiplying variables corresponding to ballot frequencies with variables corresponding to $\rho$-values or to loads). Indeed, to the best of my knowledge, there is no known profile where both PAV and MES fail core-stability simultaneously. I am also not aware of any example where the rule that maximizes the PAV score among all priceable committees (Peters and Skowron, 2020) fails core-stability.
@@ Acknowledgements
I thank Paul Gölz for useful discussions. I have used Gurobi and the cvc5 solver (Barbosa et al., 2022) in this work. This work was funded in part by the Agence Nationale de la Recherche under grant ANR22-CE26-0019 (CITIZENS) and as part of the France 2030 program under grant ANR-23-IACL-0008 (PR[AI]RIE-PSAI).

### v2

@@ 4.3 Committee size $k\geqslant 9$
The PAV-based technique that worked for up to $k=8$ does not continue to work for $k=9$, since there are examples where there is a unique global PAV committee which fails to be in the core. The following example has this property, and it is the smallest such example with respect to the number of voters ($n=27$). The indicated committee is also the unique local PAV committee.
@@ 5 Few Candidates
Inspecting the examples in Section 4 where PAV fails the core, we see that they are well-structured. Indeed, they are even laminar instances in the sense of Peters and Skowron (2020, Definition 2), and it is easy to see that on these profiles, a core-stable committee does exist. Thus, there is some hope to prove existence of core-stable committees by “patching” the PAV committee when it fails to be in the core.
We will define an artificial rule, based on PAV, and we will show that it satisfies core stability for up to $m=15$ candidates. We call it the recursive PAV rule. On a high level, the rule first computes a local PAV committee, and checks if it satisfies the core. If so, it returns it. If not, and $T$ is a deviation from $W$, it then deletes all voters who prefer $T$ to $W$, and computes a local PAV committee with respect to the remaining voters, but subject to the constraint that $T\subseteq W$. It then checks if the result is in the core; if not, it adds additional constraints until it reaches a core-stable committee. This rule is formally described using pseudocode in Algorithm 1.
For example, in the profile of Example 4.3, PAV selects the committee $W$ indicated there in blue (among other tied committees). This committee is blocked by $T=\{c_{1},c_{2},c_{3},c_{4}\}$. Thus, the recursive PAV rule would now fix all members of $T$ as winners, and maximize the PAV-score with respect to voters $v_{3}$ and $v_{4}$, obtaining the committee $W^{*}=\{c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8}\}$ (among other tied committees). This committee is in the core, so the rule terminates. Note that $W^{*}$ is not itself selected by PAV, though as we saw previously, there do exist other (global) PAV committees in the core in Example 4.3.
@@ 5.1 Analysis of the Method
where $\mathcal{A}_{t}=\{A\in\mathcal{A}:u_{A}(T_{s})\leqslant u_{A}(W_{s})\text{ for}s\in\{1,\dots,t-1\}\}$ is the set of “active” ballots. That is, $W_{t}$ locally maximizes the PAV score among all committees that include all prior deviations, taking only those voters into account that did not participate in prior deviations.
@@
              if $(W_{t},T_{t})$ is not canonical with respect to $H$ then
@@
Thus, to prove the existence of core-stable committees, it suffices to enumerate all histories and check that they fulfil the condition of Proposition 5.2. Given a potential history, one can check using an LP solver whether it is a history by checking whether the system of linear inequalities in Definition 5.1 has a solution. This way, we can compute the set of histories using a standard breadth-first search, as shown in Algorithm 2. A key insight to speed up the search is that we may break symmetries and only need to consider “canonical” histories in our enumeration. For example, we may assume without loss of generality that $W_{1}$, the first committee of the history, is $\{c_{1},\dots,c_{k}\}$. Similarly, we do not need to consider all potential deviations $T_{1}$: given our choice of $W_{1}$, the candidates in $W_{1}$ are indistinguishable to each other, as are the candidates in $C\setminus W_{1}$, and thus it suffices to take one deviation for each possible combination of the sizes of $|T_{1}\cap W_{1}|$ and of $|T_{1}\cap(C\setminus W_{1})|$. Similar symmetry-breaking conditions apply for later steps. [^6] Abstractly speaking, if an algorithm has identified sets $A_{1},\dots,A_{s}\subseteq C$ thus far (where in the present application, these sets are $W_{t}$’s and $T_{t}$’s), we can define an equivalence relation with $a\sim b$ if and only if for each $t\in[s]$, either $a,b\in A_{t}$ or $a,b\in C\setminus A_{t}$. Then two possibilities $A_{s+1}$ and $A_{s+1}^{\prime}$ for the next set in the sequence are indistinguishable if for every equivalence class of $\sim$, the two possibilities contain the same number of elements from that equivalence class. A canonical choice for the next set would choose the lexicographically first elements from each equivalence class.
@@
The recursive PAV rule fails for $m=16$, $k\in\{10,11\}$. For $m=16$, $k=10$, the smallest failure example I have found has 40 448 550 voters, though the ILP for minimizing this number did not converge within a reasonable amount of time. The example is available online. [^7] The history that witnesses this failure is the following: ($(\{c_{0},c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{9}\},\{c_{0},c_{10},c_{11}\}),$ ($(\{c_{0},c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{10},c_{11},c_{12}\},\{c_{13},c_{14},c_{15}\}),$ ($(\{c_{0},c_{1},c_{2},c_{3},c_{10},c_{11},c_{12},c_{13},c_{14},c_{15}\},\{c_{4},c_{5},c_{6},c_{7},c_{8}\})$. The recursive PAV rule does work for $m=16$, $k\in\{9,12,13,14\}$, and it is plausible that it can be fixed ad hoc for $k\in\{10,11\}$, so it is likely that the core continues to exist for $m=16$.
@@ 6 Droop Quota
On this profile, $W=\{c_{1},c_{2},c_{5},c_{6},c_{7},c_{8}\}$ is the unique global (and unique local) PAV committee for $k=6$. However $W$ is not in the Droop core: consider $T=\{c_{1},c_{2},c_{3},c_{4}\}$, which has support from $\frac{14}{24}\approx 0.583$ of the voters, while $|T|/(k+1)=\frac{4}{7}\approx 0.571$ is strictly smaller. ∎
@@ 7 Conclusions
Our approach was based on linear programming, and in particular this approach allowed us to reason independently of the number of voters. The PAV rule and its variants are particularly well-suited for these LP formulations. However, finding core counterexamples for many other rules is not possible using similar linear programs. For example, the Method of Equal Shares (MES) (Peters and Skowron, 2020) or the sequential Phragmén method (Phragmén, 1894; Janson, 2016) do not admit the same kind of linear formulations (because they would require multiplying variables corresponding to ballot frequencies with variables corresponding to $\rho$-values or to loads). The lack of such a linear formulation can be formally established using the techniques of Xia (2025). In part because computer search is difficult for these rules, to the best of my knowledge, there is no known profile where both PAV and MES fail core-stability simultaneously. I am also not aware of any example where the rule that maximizes the PAV score among all priceable committees (Peters and Skowron, 2020) fails core-stability.
@@ Acknowledgements
I thank Paul Gölz for useful discussions, and Jannik Peters and the anonymous reviewers at IJCAI 2025 for feedback that improved the presentation of the paper. I have used Gurobi and the cvc5 solver (Barbosa et al., 2022) in this work. This work was funded in part by the Agence Nationale de la Recherche under grant ANR22-CE26-0019 (CITIZENS) and as part of the France 2030 program under grant ANR-23-IACL-0008 (PR[AI]RIE-PSAI).